# Magnetic Hysteresis Modeling with Neural Operators

### v1

@@ Abstract
Hysteresis modeling is crucial to comprehend the behavior of magnetic devices, facilitating optimal designs. Hitherto, deep learning-based methods employed to model hysteresis, face challenges in generalizing to novel input magnetic fields. This paper addresses the generalization challenge by proposing neural operators for modeling constitutive laws that exhibit magnetic hysteresis by learning a mapping between magnetic fields. In particular, two prominent neural operators—deep operator network and Fourier neural operator—are employed to predict novel first-order reversal curves and minor loops, where novel means they are not used to train the model. In addition, a rate-independent Fourier neural operator is proposed to predict material responses at sampling rates different from those used during training to incorporate the rate-independent characteristics of magnetic hysteresis. The presented numerical experiments demonstrate that neural operators efficiently model magnetic hysteresis, outperforming the traditional neural recurrent methods on various metrics and generalizing to novel magnetic fields. The findings emphasize the advantages of using neural operators for modeling hysteresis under varying magnetic conditions, underscoring their importance in characterizing magnetic material based devices.
@@
Hysteresis, magnetic materials, neural operators, deep operator network (DeepONet), Fourier neural operator (FNO), rate-independent, generalization.
@@ I Introduction
Contemporary industrialization thrives on magnetic material based devices. Those devices, such as electric machines and actuators, consume more than half of the global electrical energy. Optimizing the performance of these devices includes the estimation of energy dissipated in iron as losses, which is paramount for overall performance. A considerable component of iron loss calculation is the hysteresis loss, approximated through an accurate hysteresis model [1, 2]. A material-specific hysteresis model precisely governs the constitutive relationship between the applied magnetic field ($H$) and the magnetic flux density ($B$). The relationship between these fields is nonlinear and history-dependent, traditionally modeled by the Jiles-Atherton [3] and Preisach method [4], among others [5]. Owing to the low inference cost and performance of machine learning based methods in science and engineering, deep learning based methods are frequently applied to model magnetic hysteresis, including but not limited to [6, 7, 8].
@@
This paper proposes to employ two prominent neural operators—deep operator network (DeepONet) [22] and Fourier neural operator (FNO)[23]—to model the $B$$H$ hysteresis relationship. These operator networks are trained to predict novel first-order reversal curves (FORCs) and minor loops not included in the training data. The vanilla versions of DeepONet and FNO are rate-dependent and consider the sampling rate during experiments. Further, to incorporate the rate-independent characteristic prevalent in magnetic hysteresis, this paper introduces a rate-independent Fourier neural operator (RIFNO) to predict $B$ fields when the sampling rates for the $H$ field differ between the training and test data. The primary contributions of this work are outlined as follows:
@@
  • The efficiency of neural operators is empirically demonstrated in the context of magnetic hysteresis modeling by comparing it with traditional recurrent neural architectures, highlighting their limitations. Specifically, the results are benchmarked against RNN, LSTM, and GRU models.
@@ II Motivation
This section presents the motivation to advocate employing neural operators as surrogates for modeling the hysteresis phenomenon. Hitherto, neural network based function approximation strategies have been utilized to model hysteresis. More generally, traditional neural networks, such as feedforward or recurrent neural networks, leverage the universal function approximation property [24, 25], ensuring that a continuous function $f$ can be learned to approximate the given data pairs $(\bm{x}_{i},y_{i})$ such that $y_{i}\approx f(\bm{x}_{i})$ for all $N$ data points, $1\leq i\leq N$, $i\in\mathbb{Z}$. Here, $\bm{x}_{i}$ and $y_{i}$ typically represent scalars, vectors, or tensors. The function $f(\bm{x}_{i})$ is learned as a surrogate for the underlying relationship via a neural network $\mathcal{N}(\bm{x}_{i},\theta)$, where $\theta$ represents the network parameters optimized through a gradient-based optimization method to fit the data. The learned function $f(\bm{x}_{i})$ effectively maps the finite-dimensional data $\bm{x}_{i}$ to $y_{i}$, provided the optimizer finds a suitable local minima by minimizing an appropriate loss function.
In the context of scalar magnetic hysteresis, the data pairs $(\bm{x}_{i},y_{i})$ correspond to $(H_{i},B_{i})$, representing the applied magnetic field and induced magnetic field, respectively. These data pairs are often obtained through measurements or numerical models, such as the Preisach model, which acts as a surrogate for lab experiments [6]. The index $i$ denotes the $i^{\mathrm{th}}$ measurement datum. The function $f$ encapsulates the hysteresis relationship learned from the provided data pairs. This model is expected to effectively map the finite-dimensional inputs to finite-dimensional outputs when accurately fitted. However, for novel input fields represented by a distinct $H_{i}$ sequence, the learned function $f$ fails to approximate the underlying dynamics accurately, as is presented in this work, highlighting a limitation in current strategies.
@@
In practice, it is challenging to model the hysteresis operator $\mathcal{D}$ analytically, motivating the development of neural surrogates for the operator, which are referred to as neural operators.
@@ III Methodology
Subsequently, the raw dataset is divided randomly into training and testing datasets. This results in two datasets $(h_{i},b_{i})$ and $(h_{j},b_{j})$, with $1\leq i\leq N_{\text{train}}$, $N_{\text{train}}+1\leq j\leq N_{\text{train}}+N_{\text{test}}$. Here, $N_{\text{train}}$ and $N_{\text{test}}$ represent the number of training and test samples with the relation $N_{\text{train}}+N_{\text{test}}=N_{\text{sample}}$. For the numerical experiments considered in this study, the number of training and test samples are kept equal, $N_{\text{train}}=N_{\text{test}}$, to have a fair comparison with recurrent methods. Further details regarding the number of samples are presented in Section IV. The neural operators will be trained with the dataset $(h_{i},b_{i})$ and tested with $(h_{j},b_{j})$. In the rest of the paper, we omit the dependency on time, i.e., $h_{i}$ and $b_{i}$ can both denote the functions and those functions sampled at $t_{\text{sample}}$ locations. As the fields $h_{i}$ and $h_{j}$ are distinct, the operator methods provide a framework to generalize for novel input excitation fields.
@@
While the number of hidden layers in the branch and trunk nets can vary, the number of outputs is identical. This requirement is necessary because both nets’ outputs undergo a dot product operation to approximate the $b$ fields, $\hat{b}_{i}$. Mathematically, $b_{i}\approx\hat{b}_{i}=\sum_{k=1}^{p}c_{k}\phi_{k}$, where $\phi_{k}$ represents the outputs of the branch net and $c_{k}$ represents the outputs of the trunk net. From a conventional function approximation perspective, $\phi_{k}$ can be viewed as basis functions for approximating a function, and $c_{k}$ as the coefficients of these basis functions, combined over $p$ terms to approximate the field $b_{i}$. The mean-square loss function for DeepONet is,
@@
Figure 3: Performance of neural networks on predicting the first-order reversal curves. Top row: (From Left) the input $H$ fields (in blue) for the neural network, the output $B$ fields (in black), followed by the predictions (preds.) of the $B$ fields by DeepONet, FNO, RIFNO, and LSTM, respectively. Middle row: (From Left) the reference hysteresis loop (in black) followed by the predicted hysteresis loops (in red) compared with the reference (ref.) loops (in black) for the four methods. Bottom row: (From Left) absolute errors in predicting the FORCs for the four methods (in green).
@@
DeepONet and FNO require the magnetic field’s measurement grid, the $t_{\text{sample}}$ array. However, magnetic hysteresis often exhibits the property of being rate-independent [27]. This paper proposes a minor modification in the FNO architecture to develop a rate-independent Fourier neural operator, namely RIFNO presented in Fig. 2, to incorporate the rate-independent characteristic of magnetic hysteresis. The difference between FNO and RIFNO is the input tensor $X$, which in the case of RIFNO is defined as $X=h_{i}$. Specifically, RIFNO does not concatenate the $t_{\text{sample}}$ array with $h_{i}$; hence, the architecture remains invariant to the sampling rate, making it adequate for modeling rate-independent magnetic hysteresis.
@@ IV Numerical Experiments
This section presents the numerical experiments to validate the proposed approach in predicting novel FORCs and minor loops on which training has not been performed. The three neural operator methods are compared with traditional recurrent architectures RNN, LSTM, and GRU to showcase the gain obtained with operators.
The data for the numerical experiments is generated using a Preisach-based method modeling the material NO27-1450H. The measurement data for the Preisach-based method is obtained by a Brockhaus MPG 200 D soft-magnetic material tester, using an Epstein frame calibrated to correspond with the IEC standard. The data are obtained under quasi-DC excitation, indicating that the rate of change of the magnetic flux density is controlled such that any eddy current fields are negligible, and the static hysteresis behavior is obtained. The Epstein sample strips are obtained using spark eroding and cut in the rolling direction.
@@ IV-A Error metric
IV-B First-order reversal curves
The first numerical experiment entails predicting the FORCs. This study employs the Preisach model to generate the dataset for training and testing the neural networks. The model inputs the magnetic flux density ($B$ field) and produces the corresponding magnetic field strength ($H$ field). The $b$ fields are sampled from a functional space $B$, used as the input for the Preisach model to produce the corresponding $H$ fields. This process generates data pairs ($h_{i},b_{i}$), which are used to train the neural networks. Whether one uses experimental data or any other surrogate model characterizing the material, this approach remains invariant to the choice of input sample space and output generated fields.

### v2

@@ Abstract
Hysteresis modeling is crucial to comprehend the behavior of magnetic devices, facilitating optimal designs. Hitherto, deep learning-based methods employed to model hysteresis, face challenges in generalizing to novel input magnetic fields. This paper addresses the generalization challenge by proposing neural operators for modeling constitutive laws that exhibit magnetic hysteresis by learning a mapping between magnetic fields. In particular, three neural operators—deep operator network, Fourier neural operator, and wavelet neural operator—are employed to predict novel first-order reversal curves and minor loops, where novel means they are not used to train the model. In addition, a rate-independent Fourier neural operator is proposed to predict material responses at sampling rates different from those used during training to incorporate the rate-independent characteristics of magnetic hysteresis. The presented numerical experiments demonstrate that neural operators efficiently model magnetic hysteresis, outperforming the traditional neural recurrent methods on various metrics and generalizing to novel magnetic fields. The findings emphasize the advantages of using neural operators for modeling hysteresis under varying magnetic conditions, underscoring their importance in characterizing magnetic material based devices. The codes related to this paper are at github.com/chandratue/magnetic_hysteresis_neural_operator.
@@
Hysteresis, magnetic materials, neural operators, deep operator network (DeepONet), Fourier neural operator (FNO), wavelet neural operator (WNO), rate-independent, generalization.
@@ I Introduction
Contemporary industrialization thrives on magnetic material based devices. Those devices, such as electric machines and actuators, consume more than half of the global electrical energy. Optimizing the performance of these devices includes the estimation of energy dissipated in iron as losses, which is paramount for overall performance. A considerable component of iron loss calculation is the hysteresis loss, approximated through an accurate hysteresis model [1, 2]. A material-specific hysteresis model precisely governs the constitutive relationship between the applied magnetic field ($H$) and the magnetic flux density ($B$). The relationship between these fields is nonlinear and history-dependent, traditionally modeled by the Jiles-Atherton [3] and Preisach method [4], among others [5]. Owing to the low inference cost and performance of machine learning-based methods in science and engineering, deep learning-based methods were introduced for modeling magnetic hysteresis in [6] and are frequently applied in recent works, including but not limited to [7, 8, 9].
@@
This paper proposes to employ three neural operators—deep operator network (DeepONet) [23], Fourier neural operator (FNO) [24], and wavelet neural operator (WNO) [25]—to model the $B$$H$ hysteresis relationship, which has not been done before for the modeling of magnetic hysteresis. These operator networks are trained to predict novel first-order reversal curves (FORCs) and minor loops not included in the training data. The vanilla versions of DeepONet, FNO, and WNO are rate-dependent and consider the sampling rate during experiments. Further, to incorporate the rate-independent characteristic prevalent in magnetic hysteresis, this paper introduces a rate-independent Fourier neural operator (RIFNO) to predict $B$ fields when the sampling rates for the $H$ field differ between the training and test data. The primary contributions of this work are outlined as follows:
@@
  • The efficiency of neural operators is empirically demonstrated in the context of magnetic hysteresis modeling by comparing it with traditional recurrent neural architectures, highlighting their limitations. Specifically, the results are benchmarked against RNN, LSTM, GRU, and encoder-decoder LSTM (EDLSTM) models.
@@ II Motivation
This section presents the motivation to advocate employing neural operators as surrogates for modeling the hysteresis phenomenon. Hitherto, neural network based function approximation strategies have been utilized to model hysteresis. More generally, traditional neural networks, such as feedforward or recurrent neural networks, leverage the universal function approximation property [26, 27, 28], ensuring that a continuous function $f$ can be learned to approximate the given data pairs $(\bm{x}_{i},y_{i})$ such that $y_{i}\approx f(\bm{x}_{i})$ for all $N$ data points, $1\leq i\leq N$, $i\in\mathbb{Z}$. Here, $\bm{x}_{i}$ and $y_{i}$ typically represent scalars, vectors, or tensors. The function $f(\bm{x}_{i})$ is learned as a surrogate for the underlying relationship via a neural network $\mathcal{N}(\bm{x}_{i},\theta)$, where $\theta$ represents the network parameters optimized through a gradient-based optimization method to fit the data. The learned function $f(\bm{x}_{i})$ effectively maps the finite-dimensional data $\bm{x}_{i}$ to $y_{i}$, provided the optimizer finds a suitable local minima by minimizing an appropriate loss function.
In the context of scalar magnetic hysteresis, the data pairs $(\bm{x}_{i},y_{i})$ correspond to $(H_{i},B_{i})$, representing the applied magnetic field and induced magnetic field, respectively. These data pairs are often obtained through measurements or numerical models, such as the Preisach model, which acts as a surrogate for lab experiments [7, 29]. The index $i$ denotes the $i^{\mathrm{th}}$ measurement data. The function $f$ encapsulates the hysteresis relationship learned from the provided data pairs. This model is expected to effectively map the finite-dimensional inputs to finite-dimensional outputs when accurately fitted. However, for novel input fields represented by a distinct $H_{i}$ sequence, the learned function $f$ fails to approximate the underlying dynamics accurately, as is presented in this work, highlighting a limitation in current strategies.
@@
Further, the fundamental difference between the operations of a vanilla neural network and an abstract neural operator could be described by considering the operations of a linear layer. A linear layer of a vanilla neural network maps $\bm{x}$ to $y$ through the operation, $y=w\bm{x}+\beta$, where $\bm{x}$ and $y$ represent the set of data pairs $(\bm{x}_{i},y_{i})$. Precisely, $\bm{x}\in\mathbb{R}^{m}$, $y\in\mathbb{R}^{n}$, $w\in\mathbb{R}^{n\times m}$, and $\beta\in\mathbb{R}^{n}$. However, an abstract neural operator maps the function $h(\bm{x})\in H$ to a corresponding function $b(y)\in B$ through the operation, $b(y)=K(\bm{x},y;w_{k})h(\bm{x})+\beta(\bm{x};w_{b})$, where $w_{k},w_{b}$ are the layer parameters and $K(.)$ represents the transformation kernel which differs for each operator. For instance, for FNO and WNO, the kernel takes the form of Fourier and Wavelet integrals. In contrast, for DeepONet, the kernel corresponds to the dot product operation between the branch and trunk net, described in detail in the Methodology section.
Hence, neural operators could be understood as a generalization of vanilla neural networks to approximate a higher abstract object, known as operators, mapping functions to functions. In practice, it is challenging to model the hysteresis operator $\mathcal{D}$ analytically, motivating the development of neural surrogates for the operator, which are referred to as neural operators.
Neural operators inherit the advantages of deep learning-based methods compared to the Preisach [4] and the Jiles-Atherton model [3]. One benefit of neural operators compared to the Preisach model is that it eliminates the need to rely on a specific set of measured data to characterize materials. In contrast, the classical Preisach model and its generalized versions necessitate the identification of particular branches [6]. Another benefit of deep learning-based methods is their comparatively cheap inference cost compared to the Preisach model, as discussed in several works [30, 31, 32]. Neural operators in the inference step only require a forward pass of the architecture employing algebraic and transcendental operations compared to the integral operations in the Preisach model. Also, neural operators do not need to be degaussed [33, 34], reducing the inference cost. The Jiles-Atherton model, a popular model for hysteresis modeling, faces challenges in approximating minor loops and non-symmetrical loops, as discussed in [35, 36], among others. Also, the Jiles-Atherton model requires the physical parameters before characterizing the material, which is not required for neural operators.
@@ III Methodology
The presented approach is intended to serve as a proof of concept demonstrating its effectiveness in learning and generalizing the hysteresis behavior using the Preisach model-based data [37, 38]. Modeling materials based on data obtained from models draw similarities to standard practice in machine learning-based material modeling, where model-based data are used to develop neural surrogate models. For instance, in [7, 29], among others, authors utilize the Preisach model to generate data that are further used to train the neural network-based models. This work adopts this strategy to accelerate the development and validation of neural operators.
Furthermore, neural operators have been employed for different science and engineering tasks except modeling magnetic hysteresis. The studies, including but not limited to [39, 40, 41], demonstrate that neural operators are effective even for experimental data. These advancements motivate the feasible future adoption of the proposed methodology in the magnetic domain for experimental data.
The raw dataset generated from the Preisach-based model [37, 38] for this study is divided randomly into training and testing datasets. This results in two datasets $(h_{i},b_{i})$ and $(h_{j},b_{j})$, with $1\leq i\leq N_{\text{train}}$, $N_{\text{train}}+1\leq j\leq N_{\text{train}}+N_{\text{test}}$. Here, $N_{\text{train}}$ and $N_{\text{test}}$ represent the number of training and test samples with the relation $N_{\text{train}}+N_{\text{test}}=N_{\text{sample}}$. For the numerical experiments considered in this study, the number of training and test samples are kept equal, $N_{\text{train}}=N_{\text{test}}$, to have a fair comparison with recurrent methods. Further details regarding the number of samples are presented in Section IV. The neural operators will be trained with the dataset $(h_{i},b_{i})$ and tested with $(h_{j},b_{j})$. In the rest of the paper, we omit the dependency on time, i.e., $h_{i}$ and $b_{i}$ can both denote the functions and those functions sampled at $t_{\text{sample}}$ locations. As the fields $h_{i}$ and $h_{j}$ are distinct, the operator methods provide a framework to generalize for novel input excitation fields.
@@
While the number of hidden layers in the branch and trunk nets can vary, the number of outputs is identical. This requirement is necessary because both nets’ outputs undergo a dot product operation to approximate the $b$ fields, $\hat{b}_{i}$. Mathematically, $b_{i}\approx\hat{b}_{i}=\sum_{k=1}^{p}c_{k}\phi_{k}$, where $\phi_{k}$ represents the outputs of the trunk net and $c_{k}$ represents the outputs of the branch net. From a conventional function approximation perspective, $\phi_{k}$ can be viewed as basis functions for approximating a function, and $c_{k}$ as the coefficients of these basis functions, combined over $p$ terms to approximate the field $b_{i}$. The mean-square loss function for DeepONet is,
@@
Figure 3: Neural architecture for wavelet neural operator (WNO). The input $X\!:=[h_{i},t_{\text{sample}}]$ is passed through projection tensor ($P$) and wavelet integral layers and finally downscaled ($Q$) to approximate the $B$ field.
Similar to the universal function approximation property for the neural networks [26, 27, 28], DeepONet and FNO exhibit the universal approximation property for operators as presented in [23, 19, 43, 44]. This property ensures that DeepONet and FNO can approximate any continuous operator with arbitrary accuracy, provided sufficient network parameters. However, like the universal function approximation property for neural networks, the exact number of parameters and network configuration is unknown a priori.
In the context of this work, the magnetic hysteresis phenomenon, while complex, can be modeled as a continuous operator between the applied magnetic field and the material’s magnetization response as described in Vol. I, Ch. 1 of [1]. Since the approximations required for hysteresis behavior are continuous, the universal approximation theorem provides an established mathematical foundation for employing DeepONet and FNO to learn the mapping between magnetic fields, even under complex and varying field conditions, and ensuring accurate and reliable predictions for complex field variations.
The third neural operator employed in this work is the wavelet neural operator (WNO) [25], which leverages wavelet transforms to capture the hysteresis behavior. The architecture of WNO is similar to FNO, as shown in Fig. 3. However, WNO utilizes wavelet transforms instead of Fourier transforms to learn the underlying operator.
Similar to FNO, the input of WNO is the concatenation of the input field $h_{i}$ and $t_{\text{sample}}$. The concatenated tensor, $X:=[h_{i},t_{\text{sample}}]$, is passed through a projection tensor $P$ to increase the channel size and obtain $z^{0}$ similar to FNO. Subsequently, the latent tensor $z^{0}$ undergoes a series of wavelet integral layers, parameterized by learnable weight tensors $W$ and $R$. The operation of the wavelet layers can be written as:
$z^{l+1}=\sigma\left(W^{l}z^{l}+V^{-1}(R^{l}V(z^{l}))\right),$ (4)
where $z^{l}$ represents the latent tensor at the $l^{\text{th}}$ layer, and $V$ denotes the wavelet transform, which decomposes the tensor into its wavelet coefficients. Instead of truncating certain modes (analogous to $n_{\text{m}}$ in FNO), WNO progressively decomposes the input signal into low-frequency approximations ($n_{\textrm{l}}$), to capture coarse and fine features of the signal. After applying the inverse wavelet transform, $V^{-1}$, the transformed tensor is summed with the linearly transformed latent tensor $W^{l}z^{l}$. A nonlinear activation function, $\sigma$, is applied to obtain $z^{l+1}$.
The transformations are repeated across $L$ wavelet integral layers. The output from the final layer, $z^{L}$, is passed through two projection tensors $Q$ and $\hat{Q}$ to reduce the dimensionality and produce the final predicted field $\hat{b}_{i}=\hat{Q}\sigma(Qz^{L})$. The same loss function (1), used in the case of DeepONet and FNO, is applied to approximate the field outputs.
DeepONet, FNO, and WNO require the magnetic field’s measurement grid, the $t_{\text{sample}}$ array. However, magnetic hysteresis often exhibits the property of being rate-independent [45]. This paper proposes a minor modification in the FNO architecture to develop a rate-independent Fourier neural operator, namely RIFNO presented in Fig. 2, to incorporate the rate-independent characteristic of magnetic hysteresis. The difference between FNO and RIFNO is the input tensor $X$, which in the case of RIFNO is defined as $X=h_{i}$. Specifically, RIFNO does not concatenate the $t_{\text{sample}}$ array with $h_{i}$; hence, the architecture remains invariant to the sampling rate, making it adequate for modeling rate-independent magnetic hysteresis.
The neural operator RIFNO is founded on FNO, a universal approximator for continuous operators [44]. The universal approximation property is independent of the number of input channels in the architecture. Hence, even after removing the second channel of $t_{\text{sample}}$ array, the resulting RIFNO posits to approximate the rate-independent continuous magnetic hysteresis operator.
@@ IV Numerical Experiments
This section presents the numerical experiments to validate the proposed approach in predicting novel FORCs and minor loops on which training has not been performed. The three neural operator methods are compared with traditional recurrent architectures RNN, LSTM, GRU, and recently proposed EDLSTM [14] to showcase the gain obtained with operators. The codes related to this paper are at github.com/chandratue/magnetic_hysteresis_neural_operator.
The data for the numerical experiments is generated using a Preisach-based method [37, 38] modeling the material NO27-1450H. The measurement data for the Preisach-based method is obtained by a Brockhaus MPG 200 D soft-magnetic material tester, using an Epstein frame calibrated to correspond with the IEC standard. The data are obtained under quasi-DC excitation, indicating that the rate of change of the magnetic flux density is controlled such that any eddy current fields are negligible, and the static hysteresis behavior is obtained. The Epstein sample strips are obtained using spark eroding and cut in the rolling direction.
@@ IV-A Error metric
Figure 4: Performance of neural networks on predicting the first-order reversal curves. Top row: (From Left) the predictions (preds.) of the $B$ fields by DeepONet, FNO, WNO, RIFNO, LSTM, and Encoder-Decoder LSTM (EDLSTM) respectively. Middle row: (From Left) the predicted hysteresis loops (in red) compared with the reference (ref.) loops (in black) for the six methods. Bottom row: (From Left) absolute errors in predicting the FORCs for the six methods (in green).
IV-B Hyperparameter selection
A hyperparameter search is conducted to predict the FORCs, and the obtained optimal hyperparameters are used consistently across all other cases.
For DeepONet, two key hyperparameters influencing the accuracy of the learned operator are the number of layers in the branch and trunk networks and the number of terms in the operator approximation $p$. A popular hyperparameter selection strategy, grid search [46] is performed, varying the number of layers between 2, 4, 6, and 8, and the values of $p$ between 25, 50, 75, and 100, resulting in 16 combinations. As shown in Table I, the optimal configuration, yielding the lowest error, corresponds to 8 layers and $p=25$.
For FNO, the key hyperparameters are the number of Fourier modes to retain ($n_{\mathrm{m}}$) and the number of channels for the projection $N_{\mathrm{f}}$. Similar to DeepONet, a grid search is conducted by varying $n_{\mathrm{m}}$ between 1, 2, 3, and 4, and $N_{\mathrm{f}}$ between 2, 4, 6, and 8, leading to 16 total combinations. Table II presents the results, with the best configuration identified as four modes and eight channels.
Similarly, for WNO, the important hyperparameters include the level of wavelet decomposition ($n_{\mathrm{l}}$) and the number of wavelet integral blocks $L$. A grid search is conducted by varying $n_{\mathrm{l}}$ between 1, 2, 3, and 4, and $L$ between 2, 4, 6, and 8, resulting in 16 cases. The optimal configuration, as shown in Table III, corresponded to four levels of decomposition and eight wavelet integral blocks.
The hyperparameter searches ensure the robustness of the presented results and provide optimal hyperparameter selection among the chosen cases for each neural operator.
TABLE I: Hyperparameter selection for DeepONet
Layers
$p$ Metric 2 4 6 8
25 $\mathcal{R}$ 4.96e-2 3.11e-3 1.91e-2 2.49e-3
MAE 3.19e-2 1.72e-3 1.10e-2 1.49e-3
RMSE 3.58e-2 2.25e-3 1.38e-2 1.79e-3
50 $\mathcal{R}$ 9.49e-3 4.4e-3 3.32e-3 4.34e-3
MAE 4.96e-3 2.44e-3 1.89e-3 2.34e-3
RMSE 6.86e-3 3.17e-3 2.40e-3 3.13e-3
75 $\mathcal{R}$ 2.10e-2 5.41e-3 4.93e-3 3.00e-2
MAE 1.13e-2 3.00e-3 2.71e-3 2.07e-2
RMSE 1.51e-2 3.90e-3 3.56e-3 2.17e-2
100 $\mathcal{R}$ 1.73e-2 8.91e-2 1.4e-2 5.40e-3
MAE 9.49e-3 6.01e-2 9.91e-3 2.86e-3
RMSE 1.25e-2 6.44e-2 1.05e-2 3.90e-3
TABLE II: Hyperparameter selection for FNO
Fourier modes ($n_{\textrm{m}}$)
$N_{\textrm{f}}$ Metric 1 2 3 4
2 $\mathcal{R}$ 1.51e-1 5.47e-3 8.31e-1 1.16e-2
MAE 5.37e-2 2.95e-3 5.07e-1 6.49e-3
RMSE 1.09e-1 3.95e-3 6.00e-1 8.44e-3
4 $\mathcal{R}$ 1.69e-2 4.08e-3 3.02e-3 3.84e-3
MAE 8.92e-3 2.27e-3 1.70e-3 2.12e-3
RMSE 1.22e-2 2.95e-3 2.18e-3 2.78e-3
6 $\mathcal{R}$ 1.11e-2 2.40e-3 1.96e-3 1.95e-3
MAE 6.30e-3 1.34e-3 1.13e-3 1.12e-3
RMSE 8.20e-3 1.73e-3 1.41e-3 1.41e-3
8 $\mathcal{R}$ 8.31e-3 2.25e-3 1.73e-3 1.34e-3
MAE 4.47e-3 1.28e-3 9.88e-4 7.84e-4
RMSE 6.00e-3 1.63e-3 1.25e-3 9.74e-4
TABLE III: Hyperparameter selection for WNO
Level of wavelet decomposition ($n_{\mathrm{l}}$)
$L$ Metric 1 2 3 4
2 $\mathcal{R}$ 1.12e-1 2.38e-2 2.10e-2 1.59e-2
MAE 3.21e-2 1.12e-2 9.61e-3 8.00e-3
RMSE 8.15e-2 1.72e-2 1.51e-2 1.15e-2
4 $\mathcal{R}$ 3.18e-2 6.06e-3 1.88e-2 4.67e-3
MAE 1.62e-2 3.13e-3 9.91e-3 2.46e-3
RMSE 2.29e-2 4.38e-3 1.36e-2 3.37e-3
6 $\mathcal{R}$ 1.06e-2 5.15e-3 4.01e-2 7.60e-3
MAE 5.73e-3 2.83e-3 2.22e-2 4.59e-3
RMSE 7.66e-3 3.72e-2 2.90e-2 5.49e-3
8 $\mathcal{R}$ 1.48e-2 2.10e-2 2.15e-2 2.56e-3
MAE 7.79e-3 1.28e-2 1.39e-2 1.41e-3
RMSE 1.07e-2 1.52e-2 1.55e-2 1.85e-3
IV-C First-order reversal curves
The first numerical experiment entails predicting the FORCs. This study employs the Preisach model to generate the dataset for training and testing the neural networks [37, 38]. The model inputs the magnetic flux density ($B$ field) and produces the corresponding magnetic field strength ($H$ field). The $b$ fields are sampled from a functional space $B$, used as the input for the Preisach model to produce the corresponding $H$ fields. This process generates data pairs ($h_{i},b_{i}$), which are used to train the neural networks. Whether one uses experimental data or any other surrogate model characterizing the material, this approach remains invariant to the choice of input sample space and output generated fields.